OpenStudy (anonymous):

f(x)= (x/2x+3) g(x)=2x

OpenStudy (anonymous):

oops it's backwards f(x)=x/2x+3 g(x)=2/x

OpenStudy (jiteshmeghwal9):

so wht we meed to find @cherry1239 ??